What is the magnitude of the curren t through R? DATA: R1 = R2 = RF 18.0 Ω. V1 = 6.0 Volts. V2 = 2V1. V. V. 2 2 3
Answer #1

V1 = I1R2 + I1R1 (1)

V2 = I3R3 + I2R3 (2)

I3 = I1 + I2 (3)

putting 3 into 2, we get

V2 = I1R3 + I2R3 + I2R2 (4)

solving 1 for I2 and putting into 4 and solving for I1 we get

V1 = -I2R2 + I1R1

I2 = (I1R1 - V1)/R2

I1 = (V1 + V1 + V1 x R3/R2)/(R3 + R1R3/R2 + R1)

I1 = (6 + 12 + 6x18/18)/(18 + 18x18/(18+18) + 18)

I1 = 24/(18 + 9 +18) = 0.533 A

Hence, I1 = 0.533 A